OpenStudy (anonymous):

fins the slope of the line that passes through each pair of points. (-3/4,5) & (5/4,2)

OpenStudy (anonymous):

2 - 5 = -3 5/4 + 3/4 = 8/4 -3 / 2 = 1.5